OpenStudy (anonymous):

The function g(x)-->4-3sinx is defined for the domain \[{pi /2} \le x \le A\]where A = 3pi/2. For which value of A, find the value of (g^-1)(3)

OpenStudy (experimentx):

4-3sinx = 3 solve x ..

OpenStudy (anonymous):

And then?

OpenStudy (experimentx):

what is the value of x??

OpenStudy (anonymous):

sinx= 1/3

OpenStudy (anonymous):

But then there's also inverse of g, so how do I get to the final answer?

OpenStudy (experimentx):

let x = a for now, you know for +a, sin(a) = sin(2\(\pi\) - a) and \(1\pi/2\)< \(\pi - a\) < \(3\pi/2\)

OpenStudy (anonymous):

not really following, but carry on...

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=pi+-+arcsin%281%2F3%29 answer in radians

OpenStudy (anonymous):

Thanks :)

OpenStudy (experimentx):

yw.